At Gwen's garage sale, all books were one price and all magazines were another price. Harriet bought four books and hree magazines for $1.45 and June bought two books and five magazines for $1.25. What was the price of a book and what was the price of a magazine?
Thank U:)

Answer by fractalier(6550)

Let
B = # of books and
M = # of magazines
Then we have
4B + 3M = 1.45
2B + 5M = 1.25
Then let us multiply the second equation by two and subtract them
4B + 3M = 1.45
-(4B + 10M = 2.50)
so that
-7M = -1.05
and
M = .15
Plugging this into either equation gives
4B + 3(.15) - 1.45
4B + .45 = 1.45
4B = 1.00
B = .25